The solubility of Zinc phosphate, Zn_3(PO_4)_2 is Need help
Answer #1

solubility = 5.91 *10^-4 in 100 ml = 5.91*1000/100 in one litre = 5.91*10^-3 g   = x (say)

NowKsp ( Zn3(PO4)2 =   (solubility of Zn)^3 * (solubility of PO4) ^2 = x3*x2 =x5 = (5.91*10^-3)5

                                                                   =   7.210*10^-12